A sample of 25 undergraduates reported the following dollar amounts of entertainment expenses last year: 760 745 758 774 746744 706 720 761 758 762684 697 740680747730 733 703 745 695 705 701 764 -Clickhere for the Excel Data File Round your answers to 2 decimal places.) a. Find the mean, median, and mode of this information. Mean Median Mode 732.84 $745 and 758
b. What are the range and standard deviation? Range Standard deviation 28.53 c. Use the Empirical Rule to establish an interval which includes about 95 percent of the observations interval
Answer #1

vens ann ple İ5t eltcending arrange d sample odes jnode_一 3AT-,.3.5& Ronde 4-686 angiven standard deviation is sample standard deviation. Then find 95% observations using Empirical rule.

When we have to find population standard deviation (Sigma)

Therefore use calculator then sigam = 27.95

Now

mu - 2*sigam = 732.84 - (2*27.95) = 676.94

mu + 2*sigma = 732.84 +(2*27.95) = 788.74
Answer #2

How did you come up the standard deviation?